Lita Lopez start

Help Save & Exit Submit Check my work Lita Lopez started Biz Consulting, a new business, and completed the following transactions during its first year of operations. a. Lita Lopez invested $70,000 cash and office equipment valued at $10,000 in the company. b. The company purchased an office suite for $40,000 cash. c. The company purchased office equipment for $15,000 cash. e company purchased $1,200 of office supplies and $1,700 of office equipment on credit O cash for printing an announcement of the offices opening e. The company paid a local newspaper $500 for f. The company completed a financial plan for a client and billed that client $2,800 for the service g. The company designed a financial plan for another client and immediately collected a $4,000 cash fee. Print Roferencesh. Lita Lopez withdrew $3,275 cash from the company for personal use. i. The company received $1,800 cash as partial j. The company made a partial payment of $700 cash on the equipment purchased in transaction d k. The company paid $1,800 cash for the office secretarys wages for this period payment from the client described in transaction f Required: 1. Enter the amount of each transaction on individual items of the accounting equation. (Reductions in account balances should be with a minus sign.) Liabi くPrev 1013 İİ Next >
t22 Ito yy.e c df yranna tnt d dncn 3pf yp a OMI aos e e000 dti i id rddl ve) ci ih pi ph M h hti c( t 1 1 sird ar seers r 1 a cccp i rf i esr nee 5 eceis s s t lif tuy ssa rer osoen npǐso 3 pd qh eci 3 ohs dorn as- 1 nhh 2 cahch fr afafa hf f ffh h00⑤a1h hha ci6hu ,ss 7s 733c5 au$ $$$ iii hd evedcddd The The The The 3 e the t The The The The The sh The te。 11358 25
Answer #1

As per policy, only one question is allowed to answer at a time, so answering Q1 here :

Q1) Accounting Equation:
ASSETS = LIABILITIES + EQUITY
Item Cash AR Supplies Suite Off. Equip = AP Capital Revenues Expenses
a 70000 10000 = 80000
b -40000 40000 =
c -15000 15000 =
d 1200 1700 = 2900
e -500 = -500
f 2800 = 2800
g 4000 = 4000
h -3275 = -3275
I    1800 -1800 =
j -700 = -700
k -1800                                          =                                                        -1800
Balance 14525 1000 1200 40000 26700 = 2200 76725 6800 -2300
G.Total 83425 = 83425